ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 163, 2008

Dinh, T.-C. and N. Sibony Upper bound for topological entropy of a meromorphic correspondence

29

Finston, D. R. and S. Maubach The automorphism group of certain factorial threefolds and a cancellation problem

369

Foata, D. and G.-N. Han Signed words and permutations, IV: fixed and pixed points

217

Friedman, G. Cobordism of disk knots  139

Han, G.-N. See D. Foata

Hjorth, G. Non-treeability for product group actions  383
Hollander, S. A homotopy theory for stacks  93

Izumi, M., S. Neshveyev and R. Okayasu The ratio set of the harmonic measure of a random walk on a hyperbolic group

285

Jiao, Xiaoxiang Pseudo-holomorphic curves of constant curvature in complex Grassmannians

45

Kechris, A. S. and B. D. Miller Means on equivalence relations  241
Knutson, A., E. Miller and A. Yong Tableau complexes  317

Maubach, S. See D. R. Finston

Meyerovitch, T. Tail invariant measures of the Dyck shift  61

Miller, B. D. See A. S. Kechris

Miller, E. See A. Knutson

Moretó, A. and G. Navarro Groups with three real valued irreducible characters

85

Navarro, G. See A. Moretó

Neshveyev, S. See M. Izumi

Okayasu, R. See M. Izumi

Popov, S. I. and J.-M. Strelcyn On rational integrability of Euler equations on Lie algebra so(4, ℂ)

263

Psaltis, P. A rigidity theorem for automorphism groups of trees  345

Randrianantoanina, N. Embeddings of non-commutative L^{p}-spaces into preduals of finite von Neumann algebras

1

Šemrl, P. Endomorphisms of matrix semigroups over division rings  125

Sibony, N. See T.-C. Dinh Sidi, A. A de Montessus type convergence study of a vector-valued

rational interpolation procedure

189

Strelcyn, J.-M. See S. I. Popov Yong, A. See A. Knutson

Editorial note  411